# WallFacer: Guiding Transformer Model Training Out of the Long-Context Dark Forest with N-body Problem

### v2

@@ Abstract.
In recent years, Transformer-based Large Language Models (LLMs) have garnered significant attention due to their exceptional performance across a variety of tasks. However, training these models on long sequences presents a substantial challenge in terms of efficiency and scalability. Current methods are constrained either by the number of attention heads, limiting scalability, or by excessive communication overheads. In this paper, we propose an insight that Attention Computation can be considered as a special case of n-body problem with direct interactions. Based on this concept, this paper introduces WallFacer, an efficient long-sequence training system with a novel multi-dimensional ring sequence parallelism, fostering an efficient communication paradigm and extra tuning space for communication arrangement. Through comprehensive experiments under diverse environments and model settings, we demonstrate that WallFacer significantly surpasses state-of-the-art method that supports near-infinite sequence length, achieving performance improvements of up to 77.12%.
distributed deep learning, sequence parallelism, n-body simulation, large scale training, high performance computing
copyright: noneconference: arxiv; Mar, 2025; ROTTERDAM
1. Introduction
Over the past decade, deep learning has made remarkable strides in diverse fields, including computer vision (CV) and natural language processing (NLP). Transformer-based (Vaswani et al., 2023) models, leveraging the groundbreaking attention mechanism introduced in 2018, have secured a dominant position due to their superior feature-capturing capabilities. As the technology has evolved, the ability to efficiently process long sequences has emerged as a pivotal challenge, capturing the attention of researchers and practitioners alike.
Figure 1. Comparison of the theoretical peer-to-peer communication amount for Ring Attention and different configurations of WallFacer while training a llama-7B model(Touvron et al., 2023) on 64 GPUs with Adam optimizer(Kingma and Ba, 2017) and a batch size of 4.
@@ 1. Introduction
This focus on long-sequence training and inference is underscored by its critical importance in a variety of downstream tasks. For instance, in text summarization, the ability to handle extensive sequences is vital, as the content to be summarized can range from lengthy chapters to entire books (Koh et al., 2022; Beltagy et al., 2020). Similarly, chat-based applications, such as ChatGPT (Achiam et al., 2023), require the capacity to process extensive dialogue histories to ensure conversational consistency. There are also applications in other fields like video generation(Brooks et al., 2024; Peebles and Xie, 2023) and protein structure prediction(Jumper et al., 2021; Cheng et al., 2023). In summary, the demand for efficient long-sequence training and inference mechanisms spans a broad spectrum of applications, highlighting a critical area of need within the community.
@@
The long context in the above scenarios has introduced several challenges for model training and inference: 1) Efficiency and Adaptability. The challenge of efficiency is predominantly determined by computation and communication. How to maintain high efficiency in various environments and settings has drawn great attention from academia and industry. 2) Memory. As the size of the activation during attention is quadratically related to the length of the sequence, memory consumption increases rapidly with the growth of the sequence. 3) Scalability. Large Language Models (LLMs) necessitate the use of thousands of GPUs for training, even with datasets of regular lengths. For long sequences, ensuring an acceptable scaling speedup rate is even more critical to reduce time and economic costs.
@@
Traditional parallelisms such as Data Parallelism(Hillis and Steele Jr, 1986; Shoeybi et al., 2019; Li et al., 2020; You et al., 2018), Tensor Parallelism(Shoeybi et al., 2019; Wang et al., 2022; Xu et al., 2021), and Pipeline Parallelism(Huang et al., 2018; Fan et al., 2020; Li and Hoefler, 2021; Liu et al., 2023a) have not been able to address the large memory requirement of extremely long sequences. To break through this obstacle, Sequence Parallelism has been introduced, splitting the input on the sequence length dimension. Mainstream Sequence Parallelism schemes can generally be classified into two categories: those based on all-to-all communication, and those based on ring peer-to-peer communication. Methods like DeepSpeed Ulysses(Jacobs et al., 2023), which are based on all-to-all communication, offer efficiency but require the splitting of attention heads. Consequently, these methods are limited in scalability. On the other hand, peer-to-peer communication methods(Liu et al., 2023b; Li et al., 2024), such as Ring Attention(Liu et al., 2023b), do allow for infinite context lengths; however, they necessitate the transmission of complete keys and values across all GPUs, leading to significantly high communication loads. In summary, there remains a deficiency in communication-efficient methods that are capable of supporting infinite context lengths.
In fact, we can draw valuable insights from other fields that have extensive experience in handling long sequences. In the field of n-body simulations, researchers have been dealing with vast numbers of particles, typically far exceeding the number of tokens addressed in current Transformer models. Through careful observation, we have identified that the core component of Transformers, the attention mechanism, closely resembles a special case of the n-body problem. Specifically, n-body problems with direct interactions focus on calculating the resultant force exerted by all other particles in the system on each individual particle. Similarly, in self-attention, the objective is to compute the attention score that each token in the sequence assigns to every other token. This similarity suggests that we can leverage methodologies from n-body problem research—refined over decades—to inform and enhance our approaches to tackling the relatively new challenges of long-context Transformer training.
Inspired by the methodology of n-body communication optimization(Driscoll et al., 2013) and tailored to the Transformer architecture, we introduce WallFacer, a near-infinite-context Transformer training system with multi-ring sequence parallelism that incorporates an additional parallel dimension to the existing ring-style communication, fostering an efficient communication paradigm and extra tuning space for communication arrangement. With very little extra cost of memory, WallFacer parallelism significantly reduces the peer-to-peer communication amount, as is shown in figure 1.
In summary, our paper presents these contributions:
  • We conceptualize Attention computation as a novel instance of the traditional n-body problem, providing fresh insights into optimizing and parallelizing Attention computation.
  • We offer a straightforward method that allows users to select the most suitable parallelism scheme based on their specific needs, maximizing the utility of the available tuning space within our communication framework.
  • We perform experiments on mainstream Transformer models, conducting performance and scaling tests across various computing clusters. Preliminary results indicate that our WallFacer system outperforms Ring Attention by up to 77.12%, showcasing its efficacy and scalability.
2. Background
2.1. Transformer and Long Sequence Training
The key mechanism behind Transformer-based models is Attention(Vaswani et al., 2023), which captures the text feature by calculating the attention score between every two single tokens. A standard attention function is given as:
(1) $Attention(Q,K,V)=softmax(\frac{QK^{T}}{\sqrt{d_{k}}})V$
where Q, K, and V are the Query, Value, and Key of input x with shape (B, N, H) in the case of self-attention, and $d_{k}$ represents the head dimension in commonly used multi-head attention. For convenience, we use the symbolic representations in Table 1. We can observe that in the original attention function, we need to store the intermediate value $QK^{T}$ of size $N^{2}$, which means that the memory capacity requirement and computation amount both grow in quadratic proportion with the sequence length. To solve the memory IO and capacity bottleneck, methods like flash-attention(Dao et al., 2022) have been proposed. Flash-attention is designed upon the idea of softmax decomposition (Kitaev et al., 2020; Milakov and Gimelshein, 2018; Rabe and Staats, 2022), enabling online softmax computation by storing some extra statistics. Flash-attention is thus capable of segmenting the Q, K, and V matrices into blocks, allowing the entire computation to be performed within a single kernel for each block sequentially. This strategy effectively circumvents redundant read and write operations to the High Bandwidth Memory (HBM). However, when the sequence length reaches millions, it becomes necessary to distribute the sequence across multiple GPUs. This distribution helps to reduce both the memory and computation demands on any single device. This strategy is also known as Sequence Parallelism. We will explore these methods in further detail in the following section.
Table 1. Meanings of the symbols that are used in this paper
@@ B The training batch size
2.2. Sequence Parallelism
Sequence Parallelism was first conceptualized in a study published in 2021 (Li et al., 2021), which introduced the concept as a strategy to distribute computational sequences across multiple processing steps. In this paper, we adopt a broad definition of Sequence Parallelism to include any approach that partitions a sequence during specific computational phases. Presently, Sequence Parallelism can be divided into two main categories: attention-head-sharding-based and peer-to-peer-communication-based. The former involves distributing the attention heads of multi-head attention across multiple GPUs, whereas the latter resembles a distributed version of flash-attention, relying on peer-to-peer communication to transfer keys, values, and intermediate statistics. We will now explore the advantages and disadvantages of these methods in detail.
2.2.1. Attention-head-Sharding-Based
@@
DeepSpeed Ulysses. DeepSpeed Ulysses(Jacobs et al., 2023) presents a strategy for training with long sequences by leveraging all-to-all collective communication, but is largely limited in its scalability. Functions outside the attention layer are partitioned along the sequence dimension using conventional methods. Within the attention block, Ulysses transitions from sequence parallelism to a method akin to tensor parallelism. It divides the query, key, and value matrices across the attention heads, thereby preserving the original attention computation structure. This is facilitated by two sets of all-to-all communication that switch between sequence splitting and attention head spliting.
@@
The principal merit of Ulysses lies in its simplicity and ease of implementation. Nonetheless, its reliance on the number of attention heads for partitioning activations introduces the following limitations. 1) Ulysses is limited in scalability, as it can only be expanded to as many GPUs as there are attention heads, which in turn restricts the maximum sequence length that can be processed. This limitation becomes particularly problematic when employing techniques like grouped-query attention (GQA) (Ainslie et al., 2023) or multi-query attention (MQA) (Shazeer, 2019), which further reduce the number of available attention heads for Keys and Values. 2) Ulysses encounters load-balancing issues when the number of attention heads is not evenly divisible by the number of GPUs, complicating its deployment across diverse hardware configurations. 3) The need for head-splitting makes Ulysses hard to combine with Tensor Parallelism.
@@
Megatron Sequence Parallelism. Megatron Sequence Parallelism(Korthikanti et al., 2022) shares its name with an earlier study(Li et al., 2021). Megatron Sequence Parallelism focuses on minimizing memory usage and reducing the necessity for activation recomputation. The training approach is based on Tensor Parallelism (TP). While TP is employed in the Linear and Self-Attention blocks, operations such as the LayerNorm and Dropout functions are not distributed, leading to replications within the tensor parallel group. This requires an additional all-gather operation and a further reduce-scatter operation during the forward propagation to alternate between Tensor Parallelism and Sequence Parallelism. Megatron Sequence Parallelism is not designed for training with exceedingly long sequences, as the Self-Attention and MLP layers continue to rely on Tensor Parallelism and must process the full sequence length. Like DeepSpeed Ulysses, it is also limited by the number of attention heads.
2.2.2. Ring-peer-to-peer-communication-based
@@
The primary method in peer-to-peer-communication-based strategies is Ring Attention(Liu et al., 2023b).
@@
Ring Attention. Introduced in 2023, Ring Attention(Liu et al., 2023b) innovatively partitions the sequence dimension and utilizes a ring-style peer-to-peer (P2P) communication pattern to transfer Keys and Values across all GPUs. Each GPU receives the key and value matrices from the preceding rank, updates the local attention score, and then forwards them to the next rank, as is shown in Figure 2. This method employs an online-softmax and updates attention scores incrementally, allowing the computation of attention scores without retaining the full sequence length. Thus, it potentially supports infinite context, provided sufficient computing resources are available. However, the requirement for P steps of P2P communication renders this approach less efficient in environments with high-latency communication. A more detailed discussion on this topic is presented in the subsequent section.
@@
Figure 2. An example of Ring Attention Computation on 16 GPUs in two nodes. 16 rounds of ring P2P communication are needed to reach the final results.
2.3. n-body Problem
Before diving into the model formulation, it is essential to have a brief overview of the n-body simulation. N-body simulations(Hockney and Eastwood, 1989) play a crucial role in the study and modeling of dynamics within multi-body systems. Typically, the N in n-body represents the number of entities, often reaching up to $10^{12}$, far exceeding the processing capability of a single processor. Consequently, n-body simulations are predominantly conducted on supercomputing clusters. Specifically, n-body problems involving direct interaction refer to scenarios where only the direct forces between particles, such as gravitational forces, are considered. The design of n-body simulations encounters several significant challenges similar to those discussed in previous sections: 1) Efficiency: Continuous efforts are directed towards developing more efficient frameworks and algorithms to enhance both the computational and communicational aspects of n-body simulations. 2) Memory: The number of particles in n-body problems typically surpasses those in natural language processing by orders of magnitude, presenting substantial challenges in managing memory on processors. 3) Scalability: As n-body simulations are often executed on a large amount of computing nodes, scaling the computations to more processors is notably challenging. This backdrop of challenges in n-body simulations leads us to ponder whether the decades of accumulated knowledge in this field could be leveraged to inform and guide the training of Transformer models.
3. Formulation: Transformer as N-body
Through observation, we have discerned that attention computation can be viewed as a new instance of the n-body problem with direct interactions. Moreover, the forward and backward propagation processes in neural networks can be interpreted as distinct scenarios within this n-body framework. This insight suggests that the methodologies developed for n-body problems could also be effectively applied to the long-sequence training of Transformer models.
3.1. Modeling of Attention as n-body
The general n-body problem with direct interaction can be described as:
(2) $B(x_{i})=\Sigma_{j=0}^{N}g(m_{0}(x_{j}),m_{1}(x_{i}))$
Taking the cosmology gravity problem as an example, $B(x_{i})$ represents the total gravitational force experienced by an astronomical object x within the system, $m_{0}(x)$ and $m_{1}(x)$ both denote the mass of $x$ in this case, $g$ refers to the gravitational formula between two objects, and $\Sigma$ denotes the formula for the resultant force composition. Similarly, attention computation in neural networks can be expressed in an analogous format. Moreover, this formulation allows for distinct interpretations of the processes during forward and backward propagation: Forward Propagation. From the nature of matrix multiplication and formula 1, we can tell that the forward attention computation is independent for each $Query_{i}$, and each $Query_{i}$ needs to calculate score for the $Key$s and $Value$s of every token. In this way, we can rewrite formula 1 into formula 2 by following the Flash-Attention(Dao et al., 2022) style and having:
(3) $\begin{split}m_{0}(x)&=W_{Q}x+B_{Q}=Q_{x}\\ m_{1}(x)&=W_{K}x+B_{K},\ W_{V}x+B_{V}=K_{x},V_{x}\\ g(m_{0}(x_{i}),m_{1}(x_{j}))&=softmax(\frac{Q_{xi}K^{T}_{xj}}{\sqrt{d_{k}}})V_{xj}\\ B(x_{i})&=\Sigma_{j=0}^{N}\ g(m_{0}(x_{i}),m_{1}(x_{j}))*e^{lse_{j}-lse_{new}}\end{split}$
where the Ws and Bs refer to the weights and biases in the Linear functions, while lse is the LogSumExp (Nielsen and Sun, 2016) result calculated during $g$.
Backward Propagation The primary distinction between forward and backward attention lies in the computational process during backward propagation. Specifically, to calculate the gradients for Keys and Values, it is necessary to iterate through all the Queries, and vice versa. This means that each token must interact with all three properties (Queries, Keys, Values) of every other token to compute its final gradients. We can analogize this to an n-body problem, where the interaction of forces is asymmetrical in two directions. Consequently, this requires conducting two rounds of the process as described in formula 2. In the first round, $x_{i}$ refers to $K_{i}$ and $V_{i}$ while in the second round, it refers to $Q_{i}$.
In summary, we can easily interpret both the forward and backward attention calculation into a simple form of the n-body problem.
Figure 3. An overview of the WallFacer Training System
3.2. Methodology Transfer
It is a recognized practice to harness the knowledge from traditional systems and high-performance computing to guide the design of machine learning systems. Now that we have modeled the attention calculation in Transformers as an n-body problem, we are poised to transfer a wealth of optimization methodologies from n-body simulation to Transformer training. This methodology transfer is deemed rational, as evidenced by existing techniques shared between these fields. For instance:
Particle Decomposition: Commonly used to manage memory and computation constraints, this strategy involves distributing particles across multiple processors and integrating the results using the law of force composition. This method is widely implemented in models that employ online softmax, such as Flash-Attention(Dao et al., 2022).
Hierarchical Methods: Hierarchical structures, such as the Multilevel Fast Multipole Method(Song and Chew, 1995; Lu and Chew, 1994) or tree-based processes like the Barnes–Hut simulation(Barnes and Hut, 1986), are used to approximate n-body calculations. These methods reduce the complexity from $O(n^{2})$ to $O(n\log n)$ or $O(n)$. Such approaches have been effectively applied in various fields, including computer vision, as seen in FASTERVIT(Hatamizadeh et al., 2024), etc., and in natural language processing tasks, as demonstrated in (Liu and Lapata, 2019) and so on.
Cutoff: Interactions in many systems decay with distance, making it practical to set a cutoff distance and ignore interactions beyond this threshold. This principle is similarly utilized in approximate attention algorithms like Longformer(Beltagy et al., 2020), which employs a sliding window to calculate an approximate attention score for nearby tokens while excluding those outside the window. These examples validate the feasibility of guiding Transformer system design using n-body simulation techniques. Furthermore, we have identified additional aspects of n-body simulation that could inspire innovative approaches in training long-sequence Transformer models.
In addition to the aforementioned techniques, we have explored a series of multi-dimensional algorithms used but not limited in n-body problems(Solomonik and Demmel, 2011; Aggarwal et al., 1990; Dekel et al., 1981; Agarwal et al., 1995; Gupta and Sadayappan, 1994) that optimize the communication scheme among processors while significantly reducing the total communication load, albeit at the expense of increased memory usage. The core idea involves leveraging an additional parallel dimension to partition the complete communication and computation tasks, distributing them across various devices. Inspired by these developments, we have conceived the concept of multi-ring sequence parallelism for our training system, which will be discussed in detail in the following section.
4. WallFacer Training System
4.1. Overview

### v3

# WallFacer: Harnessing Multi-dimensional Ring Parallelism for Efficient Long Sequence Model Training

@@ Abstract
Training Transformer models on long sequences in a distributed setting poses significant challenges in terms of efficiency and scalability. Current methods are either constrained by the number of attention heads or excessive communication overheads. To address this problem, we propose WallFacer, a multi-dimensional distributed training system for long sequences, fostering an efficient communication paradigm and providing additional tuning flexibility for communication arrangements. Specifically, WallFacer introduces an extra parallel dimension to substantially reduce communication volume and avoid bandwidth bottlenecks. Through comprehensive experiments across diverse hardware environments and on both Natural Language Processing (NLP) and Computer Vision (CV) tasks, we demonstrate that our approach significantly surpasses state-of-the-art methods that support near-infinite sequence lengths, achieving performance improvements of up to 77.12% on GPT-style models and up to 114.33% on DiT (Diffusion Transformer) models.
1 Introduction
@@
Over the past decade, Transformer[vaswani2023attention] models have made remarkable strides in diverse fields, including computer vision (CV) and natural language processing (NLP). As the technology has evolved, the ability to efficiently process long sequences with Transformer has emerged as a pivotal challenge. For instance, in text summarization, the ability to handle extensive sequences is vital, as the content to be summarized can range from lengthy chapters to entire books [Koh2022AnES, beltagy2020longformer]. Similarly, chat-based applications, such as ChatGPT [openai2023gpt4], require the capacity to process extensive dialogue histories to ensure conversational consistency. There are also applications in other fields like video generation[Brooks_2024, peebles2023scalable] and protein structure prediction[Jumper2021HighlyAP, cheng2023fastfold].
Figure 1: WallFacer-2 and WallFacer-4 theoretically save around 50% and 75% of total P2P communication volume for various sequence length.
@@
The long context in the above scenarios has introduced several challenges for model training and inference: 1) Efficiency and Adaptability. The challenge of efficiency predominantly lies in handling long sequences that require quadratic computations during attention, and in addressing the large amount of communication during distributed processing. 2) Memory. Besides the major obstacle of storing the model weight and optimizer states, the activation has also exceeded the capacity of a single GPU and risen as a new memory challenge due to the extreme sequence length. 3) Scalability. Current Transformer models usually require thousands of GPUs for pre-training, even with datasets of regular lengths. For longer sequences, ensuring an acceptable scaling speedup rate with both the sequence length and the number of GPUs increasing is even more critical to reducing time and economic costs.
@@
Traditional parallelisms such as Data Parallelism[hillis1986data, shoeybi2019megatron, li2020taming, you2018imagenet], Tensor Parallelism[shoeybi2019megatron, wang2022tesseract, xu2021efficient], and Pipeline Parallelism[gpipe, dapple, li2021chimera, Liu2023Hanayo] distribute the model, input batch, and the optimizer states, but can not directly address the large memory requirement of extremely long sequences as the sequence length dimension remains unchanged. To break through this obstacle, Sequence Parallelism has been introduced, splitting the input on the sequence length dimension. Mainstream Sequence Parallelism schemes can generally be classified into two categories: those based on all-to-all communication, and those based on ring peer-to-peer communication. Methods like DeepSpeed Ulysses[jacobs2023Ulysses], which are based on all-to-all communication, offer efficiency but require the splitting of attention heads. Consequently, these methods are limited in scalability and can not be scaled to more devices than the number of attention heads. On the other hand, peer-to-peer communication methods[liu2023ring, li2024distflashattn], such as Ring Attention[liu2023ring], do allow for near-infinite context lengths; however, they require the transmission of complete keys and values across all GPUs, leading to significantly high communication loads. In summary, there remains a deficiency in communication-efficient methods that are capable of supporting near-infinite context lengths.
Table 1: Meanings of the symbols that are used in this paper
@@
Inspired by the methodology of n-body communication optimization [Driscoll2013ACN], we introduce WallFacer, a novel near-infinite-context Transformer training system with multi-ring sequence parallelism that incorporates an additional parallel dimension into the existing ring-style communication. Specifically, instead of including all GPUs in a single parallel group as done in Ring Attention [liu2023ring], WallFacer groups the GPUs into teams and divides the peer-to-peer communication within these teams. This approach fosters an efficient communication paradigm and provides extra tuning flexibility for communication arrangements. With very little additional memory cost, WallFacer parallelism significantly reduces the peer-to-peer communication volume, as shown in Figure 1. Compared to previous works, WallFacer is not limited in supported sequence length by attention heads like DeepSpeed Ulysses[jacobs2023Ulysses] and Megatron Sequence Parallelism[korthikanti2022reducing], and also shows better communication efficiency and scalability than Ring Attention[liu2023ring].
In summary, our paper presents these contributions:
  • We offer a straightforward grid-search method that allows users to select the most suitable parallelism scheme based on their specific needs, maximizing the utility of the available tuning space within our communication framework.
  • We perform experiments on mainstream Transformer models, including GPT-style[gpt] and DiT-style[peebles2023scalablediffusionmodelstransformers], conducting performance and scaling tests across various computing clusters. Preliminary results indicate that our WallFacer system outperforms Ring Attention by up to 77.12% on the GPT model and up to 114.33% on the DiT model, showcasing its efficacy and scalability.
2 Background and Motivation
2.1 Transformer models and Long Sequence Training
Transformer-based models have become the mainstream approach in both Natural Language Processing (NLP) and Computer Vision (CV). GPT-style models [gpt] lead the field in NLP tasks, generating tokens in an auto-regressive manner. Meanwhile, the Diffusion Transformer (DiT) [peebles2023scalablediffusionmodelstransformers], which employs pure Transformer blocks within traditional diffusion processes, has gained widespread use over the past year for image and video generation. The key mechanism behind these Transformer-based models is attention[vaswani2023attention], which captures the text feature by calculating the attention score between every two single tokens. A standard attention function is given as:
$Attention(Q,K,V)=softmax(\frac{QK^{T}}{\sqrt{d_{k}}})V$ (1)
where $Q$, $K$, and $V$ are the query, value, and key of input x with shape ($B$, $N$, $H$) in the case of self-attention, and $d_{k}$ represents the head dimension in commonly used multi-head attention. For convenience, we use the symbolic representations in Table 1. We can observe that in the original attention function, we need to store the intermediate value $QK^{T}$ of size $N^{2}$, which means that the memory capacity requirement and computation amount both grow in quadratic proportion with the sequence length. To solve the memory IO and capacity bottleneck, methods like flash-attention[dao2022flashattention] have been proposed. Flash-attention is designed upon the idea of softmax decomposition [kitaev2020reformer, milakov2018online, rabe2022selfattention], enabling online softmax computation by storing some extra statistics. Flash-attention is thus capable of segmenting the $Q$, $K$, and $V$ matrices into blocks, allowing the entire computation to be performed within a single kernel for each block sequentially. This strategy effectively circumvents redundant read and write operations to the High Bandwidth Memory (HBM).
However, the sequence length can reach hundreds of thousands, when dealing with multi-round chatting, or high-resolution long video generation. It then becomes necessary to distribute the sequence across multiple GPUs. This distribution helps to reduce both the memory and computation demands on any single device. This strategy is also known as sequence parallelism. We will explore these methods in further detail in the following section.
2.2 Sequence Parallelism
Sequence parallelism was first conceptualized in a study published in 2021 [Li2021SequencePL], which introduced the concept as a strategy to distribute computational sequences across multiple processing steps. Presently, Sequence parallelism can be divided into two main categories: attention-head-sharding-based and peer-to-peer-communication-based. The former involves distributing the attention heads of multi-head attention across multiple GPUs, whereas the latter resembles a distributed version of flash-attention, relying on peer-to-peer communication to transfer keys, values, and intermediate statistics.
2.2.1 Attention-Head-Sharding-Based
@@
DeepSpeed Ulysses. DeepSpeed Ulysses[jacobs2023Ulysses] presents a strategy for training with long sequences by leveraging all-to-all collective communication, but is largely limited in its scalability. Within the attention block, Ulysses transitions from sequence parallelism to a method akin to tensor parallelism with two all-to-all communication. It divides the query, key, and value matrices across the attention heads, thereby preserving the original attention computation structure.
@@
The principal merit of Ulysses lies in its simplicity and ease of implementation. Nonetheless, its reliance on the number of attention heads for partitioning activations introduces the following limitations. 1) Ulysses is limited in scalability, as it can only be expanded to as many GPUs as there are attention heads, which restricts the maximum sequence length that can be processed. This limitation becomes particularly problematic when employing techniques like grouped-query attention (GQA) [ainslie2023gqa] or multi-query attention (MQA) [shazeer2019fast], which further reduce the number of available attention heads for Keys and Values. 2) Ulysses encounters load-balancing issues when the number of attention heads is not evenly divisible by the number of GPUs, complicating its deployment across diverse hardware configurations. 3) The need for head-splitting makes Ulysses hard to combine with Tensor Parallelism.
@@
Megatron Sequence Parallelism. Megatron Sequence Parallelism focuses on minimizing memory usage and reducing the necessity for activation recomputation. The training approach is based on Tensor Parallelism (TP). While TP is employed in the Linear and Self-Attention blocks, operations such as the LayerNorm and Dropout functions are not distributed, leading to replications within the tensor parallel group. This requires an additional all-gather operation and a further reduce-scatter operation during the forward propagation to alternate between Tensor Parallelism and Sequence Parallelism. The Self-Attention and MLP layers continue to rely on Tensor Parallelism and must process the full sequence length. Like DeepSpeed Ulysses, it is also limited by the number of attention heads.
2.2.2 Ring-peer-to-peer-communication-based
@@
The primary method in peer-to-peer-communication-based strategies is Ring Attention[liu2023ring].
@@
Ring Attention. Introduced in 2023, Ring Attention[liu2023ring] innovatively partitions the sequence dimension and utilizes a ring-style peer-to-peer (P2P) communication pattern to transfer Keys and Values across all GPUs. Each GPU receives the key and value matrices from the preceding rank, updates the local attention score, and then forwards them to the next rank, as is shown in Figure 2. This method employs an online-softmax and updates attention scores incrementally, allowing the computation of attention scores without retaining the full sequence length. Thus, it potentially supports infinite context, provided sufficient computing resources are available. However, the requirement for $P$ steps of P2P communication renders this approach less efficient in environments with high-latency communication.
@@
Figure 2: An example of Ring Attention Computation on 16 GPUs in two nodes. The Communication is largely limited by the inter-node bottleneck.
Figure 3: An overview of the WallFacer Training System
2.3 Motivation
Through observation, we identify two main drawbacks of Ring Attention. First, the communication overhead is exceedingly high because every GPU in the system must send and receive keys and values for nearly the entire sequence length before completing the attention computation. Second, variations in bandwidth between and within computing nodes can cause communication bottlenecks. As illustrated in Figure 2, the bandwidths between GPUs 3 and 4 and between GPUs 11 and 12 are lower than those between other GPUs in the ring. Despite this, the system is forced to operate in a complete circle, which can result in unnecessary idle times for other GPUs.
Inspired by techniques used in n-body simulations [Hockney_Eastwood_1989], we propose an approach that addresses both challenges simultaneously. In n-body simulations, such as those calculating direct forces between particles (e.g., gravitational forces), researchers discovered that communication overhead can be minimized by introducing additional dimensions of communication [Driscoll2013ACN]. Similarly, we propose grouping GPUs into teams within the ring-style sequence parallelism framework and distributing communication tasks among team members. This approach reduces the overall communication volume and mitigates the bandwidth bottlenecks previously mentioned. Building upon this concept, we developed the WallFacer training system, which we will detail in the following section.
3 WallFacer Training System
3.1 Overview